Use Hess’s law to calculate ∆H° for the reaction:

C(s) + 2H2(g) + ½O2(g) → CH3OH(l)   ∘= ?

using only the following data:

H2(g) + ½O2(g) → H2O(l)                   = -285.8 kJ

C(s) + O2(g) → CO2(g)                    = -393.5 kJ

2CH3OH(g) + 3O2(g) → 2CO2(g) + 4H2O(l)                = -1452.8 kJ  
Answer #1

The given reactions are as follows :

(1) H2(g) + ½O2(g) → H2O(l) ∆H° = -285.8 kJ

(2) C(s) + O2(g) → CO2(g)        ∆H° = -393.5 kJ

(3) 2CH3OH(g) + 3O2(g) → 2CO2(g) + 4H2O(l) ∆H° = -1452.8 kJ

For required reaction: (2) + 2(1) - 0.5(3)

Then required delta H is as follows :

Delta H = -393.5 + 2(-285.8) - 0.5(-1452.8)

Delta H = 238.7 KJ

Therefore, answer is 238.7 KJ.

Please give good rating.